OpenStudy (anonymous):

At a certain point in time, each dimension of a cube of ice is 30 cm and is decreasing at the rate of 2 cm/hr, so how fast is the ice melting (losing volume)?

jigglypuff314 (jigglypuff314):

Volume of cube = (side)^3 so V = s^3 then take the derivative of that :)

OpenStudy (anonymous):

3s^2

OpenStudy (campbell_st):

this is related rates \[\frac{dv}{dt} = \frac{dv}{dl} \times \frac{dl}{dt}\]

jigglypuff314 (jigglypuff314):

don't forget chain rule! :)

OpenStudy (anonymous):

so, 6s?

jigglypuff314 (jigglypuff314):

not quite :) I meant that the derivative of V = s^3 should have been V' = 3s^2 (s')

OpenStudy (campbell_st):

l is the side length and you know dl/dt = 2 you know V = l^3 find dV/dl by differentiating volume... then subtitute your given value of l

OpenStudy (anonymous):

so, after v'=3s^2, i do?

jigglypuff314 (jigglypuff314):

V' = 3s^2 (s') you were given s' = -2 and that s = 30 so plug that in and you'll get V' = 3 (30)^2 (-2) =

OpenStudy (anonymous):

-5400

jigglypuff314 (jigglypuff314):

yup :)

OpenStudy (anonymous):

<3